Infuse heparin 20,000 units in 500 ml of normal saline IV over 12 hours. The administration set has a drop factor of 15 gtt/ml. The nurse will set the gtt per minute at what rate?
Answer #1

500 ml has to be infused in 12 hours.

Drip rate = 500 /12 ml/hour

Drop factor = 15 gtt/ml

gtt per hour = 500/12 * 15= 625 gtt/hour

gtt per minute = 625/60 = 10.4 gtt/min